OpenStudy (anonymous):

Point E is located at (2, −3), and point F is located at (−2, −1). Find the y value for the point that is 3 over 4 the distance from point E to point F. 1.75 0.5 −2 −1.5

OpenStudy (anonymous):

Midpoint Pt(x,y): \[\frac{ x_{1}+x_{2} }{ 2},\frac{ y_{1}+y_{2} }{ 2}\] (2, -3) (-2, -1) Mpt = (0, -2) (0, -2) (-2,-1) Mpt = (-1,-1.5) 3/4 Distance from E to F so... it's -1.5

OpenStudy (anonymous):

appreciate it